Consider the 300 households surveyed in the DASCO Electronics example that actually
purchased a large screen Television. These respondents were asked to identity how the
purchased was transacted
Type of Purchase No. of Respondent
In store 183
Internet 87
Mail Order 30

Calculate the probability of picking a household that purchased the television via Internet or
by Mail order if the one respondent is randomly selected from a sample of 300 household.
1
Expert's answer
2020-06-24T17:25:03-0400

Formula: p( A or B) = p(A) +p(B)


Solution: p( via internet or mail)= p(internet)+p( mail)

= 87/300 + 30/300

= 117/300=0.39